### v1

# AutoGraphAD: A novel approach using Variational Graph Autoencoders for anomalous network flow detection

Abstract
Network Intrusion Detection Systems (NIDS) are essential tools for detecting network attacks and intrusions. While extensive research has explored the use of supervised Machine Learning for attack detection and characterisation, these methods require accurately labelled datasets, which are very costly to obtain. Moreover, existing public datasets have limited and/or outdated attacks, and many of them suffer from mislabelled data. To reduce the reliance on labelled data, we propose AutoGraphAD, a novel unsupervised anomaly detection approach based on a Heterogeneous Variational Graph Autoencoder. AutoGraphAD operates on heterogeneous graphs, made from connection and IP nodes that capture network activity within a time window. The model is trained using unsupervised and contrastive learning, without relying on any labelled data. The reconstruction, structural loss, and KL divergence are then weighted and combined in an anomaly score that is then used for anomaly detection. Overall, AutoGraphAD yields the same, and in some cases better, results than previous unsupervised approaches, such as Anomal-E, but without requiring costly downstream anomaly detectors. As a result, AutoGraphAD achieves around 1.18 orders of magnitude faster training and 1.03 orders of magnitude faster inference, which represents a significant advantage for operational deployment.
1 Introduction
In recent years, attacks and intrusions have been a growing problem. The attacks reported each year have been growing exponentially, with attacks becoming more and more sophisticated [9623451]. Thus, more robust Network Intrusion Detection Systems (NIDS) are needed to deal with the increasing volume and complexity of network attacks.
A substantial body of research has explored the use of Machine Learning (ML) to detect and characterise attacks [8171733, LightweightIDS, 10127959, 10.1007/978-981-13-1921-1_35]. However, most of these approaches are susceptible to adversarial attacks, and therefore do not suit real-life deployment [7467366]. Graph Neural Networks are known to be more robust against adversarial attacks, since they do not rely solely on network data, but also learn the structures and relationships among network traffic flows [pujolperich2021unveilingpotentialgraphneural, 10123384, 10.1145/3664476.3664515].
In the NIDS domain, the number of datasets that can be used to create and train ML models remains limited [goldschmidt2025networkintrusiondatasetssurvey]. Most of these datasets use outdated attacks and are highly synthetic, creating a distorted image of how attacks are structured and work [PINTO2025111177].
Labelling datasets is a costly and laborious task, which explains the scarcity of high-quality labelled data sets [Braun_2024]. The main challenge is that each network flow must be examined and annotated by a domain expert, such as a Security Analyst [GUERRA2022102810]. This makes labelling large real-world network data almost impossible, which is why most research in this area relies on synthetic datasets [GUERRA2022102810].
Anomaly detection is a good approach to mitigate reliance on labelled datasets, as we only need normal data to establish a baseline, and anything that deviates from it is considered anomalous. In the domain of graph-based anomaly detection for NIDS, Anomal-E is arguably the most representative and influential approach [Caville_2022]. Anomal-E constructs a homogeneous graph, where network flows are represented as edge features, which are then transformed to edge embeddings for anomaly detection. However, Anomal-E requires the generation of large embeddings, introducing significant overhead and slowing down the detection process [doi:10.1126/science.153.3731.34]. Furthermore, in the case of concept drift [Webb2016-vo], the entire pipeline has to be re-trained, making it computationally expensive to deploy in operational environments.
In the case of Anomal-E, and many other approaches [arganids, zoubir2024integratinggraphneuralnetworks, 10.1007/978-3-030-50423-6_12], they rely on traditional downstream anomaly detection algorithms, which are often not suitable for a streaming environment [app13106353], which makes them very difficult to deploy in real-world networks.
In this paper, we propose AutogGraphAD, a novel approach for unsupervised anomaly detection based on Variational Graph Autoencoders (VGAEs). In our approach, we represent network traffic as a heterogeneous graph with two distinct types of nodes, IP and Connection nodes. We then performed unsupervised contrastive training for the VGAE, without requiring any labelled data. We leverage the generated embeddings to calculate anomaly scores that are then used for anomaly detection. We have tested multiple different architectural variations to determine which implementation yields the best results at different levels of contamination (i.e., presence of anomalous flows) of the training dataset. AutoGraphAD yields the same, and in some cases better, results than Anomal-E while not relying on the use of costly downstream estimators. Thus, faster training by 1.18 orders of magnitude and faster inference by 1.03 orders of magnitude is achieved, showcasing a significant advantage for operational deployment.
2 Problem Description
As mentioned in Section 1, the main problem that this paper aims to solve is the detection of network attacks without the need for expensive and limited labelled data. This is a very important problem that needs to be solved if we want to achieve more efficient results and a better real-life deployment of Machine Learning based NIDS.
The main issue with current solutions such as Anomal-E [Caville_2022] and ARGANIDS [arganids] is the reliance on downstream detectors such as HBOS [goldstein2012hbos] to classify the embeddings generated for classification purposes. This adds an additional layer of computation that needs to be retuned often to account for concept drift [Webb2016-vo] and the network’s dynamic nature [app13106353].
Moreover, certain one class classification algorithms tend to become prohibitively expensive when depending on the size of the embedding used. This can be seen in the case of OneClassSVM [NIPS1999_8725fb77] and its linear variants, where the dimensionality of the features can negatively impact training and inference time, as well as memory complexity and overall performance [scikit-learn]. It should be noted that, in the case of Anomal-E [Caville_2022], the reported results use 256 vector embeddings. These embeddings can be considered large and affect anomaly detection algorithms as they suffer from the curse of dimensionality [doi:10.1126/science.153.3731.34].
An interesting approach to avoid reliance on a downstream estimator is the use of dynamic Graph Autoencoders (VGAEs) to detect anomalies. This approach has been used successfully in network monitoring as seen in GAT-AD [LATIFMARTINEZ2025110830]. In the case of GAT-AD, the predicted and actual values are used to calculate an anomaly score, and then, based on those node scores, we can detect which nodes are anomalous and which are not.
In the case of anomaly detection in a network setting, the majority of state-of-the-art approaches rely on downstream estimators to conduct the anomaly detection.
In this paper, our aim is to achieve network detection without relying on a downstream detector. We rely solely on reconstruction errors and metrics and not on any other downstream detectors. We are aiming to leverage VGAEs and their reconstruction errors as a way to detect anomalous nodes. AutoGraphAD is substantially faster computationally than Anomal-E’s and ARGANID’s approaches, while much simpler and much more easily tunable.
3 Background
In this section, we describe the main pre-requisites of the technologies that we will be using in this paper as part of developing our solution.
3.1 Variational Graph Autoencoders
The underlying technology that we will use in this paper is VGAEs. VGAEs were originally proposed by Kipf in 2016 [kipf2016variationalgraphautoencoders].
They are a Graph Neural Network extension of traditional Variational Autoencoders [kingma2022autoencodingvariationalbayes]. Their initial design relied on the reconstruction of the graph structure combined with the loss of Kullback-Leibler regularisation [kipf2016variationalgraphautoencoders].
The variational autoencoder works by the encoder component creating uncertainty-aware node embeddings that are then parametrised into node embeddings and coupled with a Dot Product to recreate the Adjacency matrix of the graph [kipf2016variationalgraphautoencoders]. Then both adjacency matrices are used to compute the reconstruction error. The equations used to calculate the errors can be seen in the following: Equation 1 is the Dot Product, Equation 2 is the Binary Cross Entropy Loss used to compute the structural reconstruction error [kipf2016variationalgraphautoencoders], of Equation 3 is the KL regularisation error, and Equation 4 is the total loss used.
$A^{\prime}=Z\cdot Z^{T}$ (1)
$L_{Recon}=BCE(A^{\prime},A)$ (2)
$KL_{Reg}=-\frac{1}{2}*\sum(1+\log(\sigma^{2})-\mu^{2}-\sigma^{2})$ (3)
$L_{Total}=L_{Rrecon}+KL_{Reg}$ (4)
3.2 GraphSAGE
GraphSAGE is a Graph Neural Network architecture that can learn to generate embeddings from sampling and aggregating the node’s neighbours [NIPS2017_5dd9db5e]. After the aggregation, the generated embedding goes through a non-linear update.
The main advantage of GraphSAGE compared to other algorithms is that it can easily work with unseen nodes [NIPS2017_5dd9db5e], making it ideal for real-time deployment where the nodes and graph structure differ from training data. In addition, GraphSAGE has been proven to scale very well in larger graphs [NIPS2017_5dd9db5e], making it an ideal choice for a large network deployment.
3.3 Adversarially Regularised VGAEs
Adversarially Regularised VGAEs [pan2019adversariallyregularizedgraphautoencoder] is another approach to train the autoencoders to generate higher quality embeddings. This approach uses the exact same approach as the original VGAE paper [kipf2016variationalgraphautoencoders] but adds the use of a regulariser model to improve the performance. The regulariser works by taking samples from the generated embeddings and from the a Gaussian distribution and learns to separate which one is which. This is then fed into the total loss that is used for back-propagation.
3.4 Multiple Sampling Variational Autoencoders
Another upgrade that can be done to enhance the VGAEs is through the use of multiple latent embeddings per node. The idea behind using multiple embeddings that are then used to reconstruct multiple reconstruction losses. This allows for the generation of an unbiased estimate of the same objective with a much smaller variance [pmlr-v48-mnihb16]. This has been proven to improve the performance of the model and yield better results [pmlr-v80-buchholz18a].
3.5 KL Annealing
KL Annealing can be used to improve the performance of VAEs [bowman2016generating]. KL Annealing works by slowly increasing the KL regularisation loss through a set number of epochs. By slowly increasing the effect of the KL loss, the autoencoder is initially forced to learn meaningful embeddings, and then transitions on enforcing prior matching. The goal of this approach is to mitigate the posterior collapse, where KL transitions to 0, which can negatively affect the performance of the VAE.
3.6 Graph Masked Autoencoders
Masking was proposed by Hou et al. [GraphMAE] in the Graph Masked Autoencoder (GraphMAE) and used in the Heterogenous Graph Masked Autoencoder (HGMAE) [tian2023heterogeneousgraphmaskedautoencoders] to enhance the performance of their model. Their GAE architecture is enhanced by masking the input node features as well as the latent embeddings. Additionally, as part of model training, the graph edges are dropped. For masking, they use a learnable masking token similar to the one used in BERT [devlin2019bertpretrainingdeepbidirectional]. Masking and edge dropping make it harder for the autoencoder to reconstruct the information, and thus the autoencoder does not learn trivial solutions during training but rather focusses on learning the node relations.
4 Methodology
In this section, we focus on describing our methodology and how we have created, trained, and evaluated AutoGraphAD. Additionally, we describe our pre-processing, graph-building, and dataset creation pipelines.
4.1 Dataset Used
For this work, we have chosen to use the UNSW-NB15 dataset [7348942]. We chose UNSW-NB15 for the AutoGraphAD evaluation, as it is a widely used dataset with clearly labelled modern attacks. Thus, our evaluation is better suited for the current environment. The data set has been generated from the raw PCAP files using NFStream [AOUINI2022108719]. For the generation of the flows, we used the same settings as were used in the original dataset generation. NFStream can output flows with multiple features. We chose the features that overlap with the features present in NetFlow V9 [claise2004cisco].
For UNSW-NB15, we used the generated labels from the original paper [7348942] and matched them with the flows generated by NFStream. The features used can be found in Table V.
4.2 Dataset Creation
Having described the data pre-processing, it is important to describe how the datasets are created. The goal of this paper is to explore and showcase our model’s ability to detect anomalies even when the training data are contaminated. To achieve this, we have created different training parts with different levels of contamination similar to Anomal-E [Caville_2022]. The main difference between Anomal-E’s approach and ours is that Anomal-E uses 10% of the data set to create a single large graph for the training, and the same is repeated for testing. Our goal is to create 3 levels of contamination: No contamination, 3.5% Contamination, and 5.7% Contamination.
In our case, we use time windows to detect attacks and simulate a real-life deployment. For the training split of the dataset, we follow the following procedure; once we create all the sliding windows for the entire dataset, we calculate if the windows are anomalous or not, meaning if they contain attacks or not, and we do a stratified splitting to all datasets. For a training dataset that does not have any attacks, we removed all the anomalous graphs. For 3.5% contamination, we keep the stratified split, and for 5.7% we downsample the benign graphs for UNSW-NB15 as the normal contamination of the level is around 3.5%.
4.3 Feature Pre-processing and Graph Creation
For feature pre-processing, we use a similar pre-processing approach to the one proposed in Anomal-E [Caville_2022]. The dataset is initially checked for flows with missing or infinite values, which are then removed. Following this step, we begin by building the time windows that will be used for training, testing, and evaluation. We then split the randomly generated graphs into parts for training, evaluation, and testing. We allocate 70% of the data for training, 20% for evaluation, and 10% for testing. For the time windows, we chose to create non-overlapping graphs that capture 180 seconds of traffic.
The second step of the pre-processing is feature normalisation. For categorical features, which in our case is the IP version and protocols, we employ One-Hot encoding [Hancock2020]. In the case of the protocols, we use Rare Labelling from the library Feature Engine [Galli2021] due to the incredibly large number of protocols and the fact that the majority of these are very sparse in the data. Including these features in One-Hot encoding would increase the dimensionality of the data without yielding any performance gain; as such, the rarer protocols are bundled in a third category labelled Rare, resulting in three One-Hot categories, two for the two most common features, and one for all the rare features.
For the continuous features, we are using L2 row wise normalisation. This choice was made because it does not require any significant pre-processing and can be done on the fly in a real-life environment, allowing for an easier real-life deployment. The entire process from dataset creation to graph creation can be seen in Figure 1.
Figure 1: Dataset Pre-processing pipeline for graph generation
4.4 Network Architecture and Variants
Figure 2: The proposed Architecture of AutoGraphAD in a training setting. AutoGraphAD mainly focuses on the reconstruction and use of the connection nodes as the IP nodes have placeholder values. The pipeline starts with the encoder that generates the latent space embeddings that are then reparametised to create the embeddings that are then used for structure and feature reconstruction. Using the reconstructed values, we calculate the reconstruction losses used for back-propagation. The GNN encoder and decoder can be switched out with every other GNN algorithm. The same can be done regarding the losses that are used in Back-propagation.
Our proposed architecture builds upon the original VGAE [kipf2016variationalgraphautoencoders] and extends it by incorporating elements from GraphMAE [GraphMAE] and HGMAE [tian2023heterogeneousgraphmaskedautoencoders]. Our architecture is designed to reconstruct both node features and graph structure during training. The main upgrades introduced in our design are as follows.
We are using the same masking component as used in GraphMAE [GraphMAE] and HGMAE [tian2023heterogeneousgraphmaskedautoencoders], but instead of masking both latent embeddings and original graph embeddings, we only mask the original graph nodes. We made this decision because during the evaluation of AutoGraphAD there was a negligible improvement when using double masking.
We incorporated random edge drop during training to force the model to learn non-trivial relationships between nodes [tian2023heterogeneousgraphmaskedautoencoders]. We have also incorporated negative edge sampling as a contrastive learning objective to help the autoencoder to better distinguish between real and fake edges. This helps our autoencoder to learn better the structure of the graphs.
For structural reconstruction, we use an enhanced Dot Product. The enhanced Dot Product uses learnable weights to better capture the structural dependencies of the graphs. This is because a heterogeneous graph can be split into smaller bipartite graphs [Shi2022]. This allows us to assign a learnable weight for each of the edge types and better capture that individual relationship. For the structural reconstruction error, we use Binary Cross entropy loss as seen in the original VGAE paper [kipf2016variationalgraphautoencoders].
$A^{\prime}=Z\cdot(W*Z^{T})$ (5)
Regarding feature reconstruction, a GNN has been employed to decode the node features. A GNN was chosen as it is the most capable of decoding the embeddings correctly. GNNs as a decoder are used in both HGMAE and GraphMAE [tian2023heterogeneousgraphmaskedautoencoders, GraphMAE], but in these approaches they have focused only on decoding masked embeddings. In this architecture, all nodes are decoded. In our approach, we only decode the connection nodes as the IP nodes are placeholder values that do not contain any meaningful information.
Regarding the feature reconstruction errors, we have created two main variants, one using the Cosine Embedding Error and one using the Mean Squared Error (MSE). The Cosine Embedding Error approach has been inspired by HGMAE [tian2023heterogeneousgraphmaskedautoencoders] and GraphMAE [GraphMAE], while the MSE approach has been chosen as the goal here of the autoencoder is to be able to generate meaningful embeddings in which the decoded values must be closest to the actual values. The KL divergence loss used in our model is the same as in the original VGAE paper [kipf2016variationalgraphautoencoders].
For the total reconstruction loss, we combine the feature reconstruction loss with the structural reconstruction loss and KL divergence. The total loss is then back-propagated through the model to update the model. Each component of the total reconstruction loss, with the exception of the KL divergence, is weighted by a user defined value. The model architecture can be found in Figure 2 and the total loss is depicted in equation 6.
$L_{Total}=\alpha*L_{Structure}+\beta*L_{Features}+KL$ (6)
4.5 Model Training
The model was trained with 100 maximum allowable epochs. We used an early stopping mechanism during the training to prevent overfitting and expenditure of the computational resources. Furthermore, we saved the model’s weights at the point that the model achieved its best performance. We ensured that we maintain the optimal version of the model.
4.6 Anomaly Detection
For anomaly detection, we are calculating the anomaly scores of the connection nodes. The anomaly scores are based on the calculation of the reconstruction errors with the KL term. The idea is that the anomaly score will be lower for normal nodes and higher for anomalous nodes. To determine the threshold of what an anomaly is and what it is not, we use percentiles to determine a user defined percentile, and then everything above it is determined as anomalous scores. For our score, we use the following equation.
$Score=\alpha*L_{Feat}+\beta*L_{Struct}+\gamma*KL$ (7)
The equation works by using a user-assigned score to weight the importance of each loss in the anomaly score. This allows us to use different weights for different settings. Additionally, we employ Robust Scaling [https://doi.org/10.1002/widm.1236] of all reconstruction errors and KL Divergence that are used for the anomaly scores. This guarantees that all scores are on the same scale and that no value dominates the range and skews the anomaly scores [10.1145/2594473.2594476]. Our entire approach runs in PyTorch and on the GPU. Providing a faster processing time that using Sci-kit learn [scikit-learn]. The pipeline can be found in Figure 9 in the Appendix.
4.7 Evaluating Inference Performance
To evaluate the performance of each method, we have chosen to measure the training and inference times. For timing, we used the Python built-in library, time. More specifically, we used the function perf_counter_ns() which gives us the most accurate measurement for our experiment, since we are measuring short durations.
For timing, we do not count the parts where data transfer takes place to make the comparisons more equal. Moreover, in the case of AutoGraphAD, we can time each individual pass when generating the Anomaly Score, but this is not possible during threshold setting and threshold inference timing and when timing Anomal-E’s downstream estimator training and inference. For these cases, we get the total time and divide it based on the number of graphs that were processed.
For the recorded timings, in the case of AutoGraphAD we add the average time per graph during the anomaly score with its respective time in the threshold component. For example, in the case of the training time, we add the average anomaly score generation time per graph with the estimated threshold calculation time per graph.
4.8 Libraries Used
To create our VGAE, we used PyTorch [NEURIPS2019_bdbca288] and an extension of PyTorch called PyTorch Geometric [fey2019fastgraphrepresentationlearning] which is best suited for the creation and training of GNNs. To ensure easier reproducibility and faster training, we employed PyTorch Lightning [Falcon_PyTorch_Lightning_2019]. Regarding anomaly detection, we employed NumPy [harris2020array] and PyOD [zhao2019pyod]. Finally, for data pre-processing, window creation, and running the downstream tasks for Anomal-E, we used Scikit-learn [scikit-learn] and Pandas [mckinney-proc-scipy-2010]. To run the Anomal-E code we used DGL [wang2020deepgraphlibrarygraphcentric].
5 Experimental Results
5.1 Experimental setup
To evaluate AutoGraphAD, we used UNSW-NB15 in both AutoGraphAD and Anomal-E. For Anomal-E, we used an implementation found in a GitHub repository [^1] https://github.com/waimorris/Anomal-E.
Overall, we trained the models in the different contamination levels [^2] The contamination level indicates the percentage of anomalous flows present in the training set. Ideally, anomaly detection methods should be trained only on normal flows. However, this is difficult to guarantee in real-world environments. Hence, robustness to a small degree of contamination is desirable for practical deployment in operational networks. followed by the anomaly detection estimators’ grid search for Anomal-E and weight search for the Anomaly Score in the VGAE implementation. During the hyperparameter search, the model weights were frozen.
To run our experiments, we used a server with an Nvidia RTX 3090 with 24 GB of VRAM. The CPU of the server is an AMD Ryzen 3950X, 16-Core Processor. The server came with 64 GB of RAM and its operating system was Ubuntu 22.04.4 LTS.
5.2 VGAE Variants
We exhaustively tested multiple different VGAE variants. Our base architecture is node masking and edge dropping to force the variational autoencoders to learn non-trivial node relationships. The variants that we evaluated are the following: $(1)$ Base architecture using MSE for node feature reconstruction, $(2)$ Base architecture using Cosine Embedding Error for node feature reconstruction, $(3)$ Regularised Base Architecture, $(4)$ Split Base Architecture, and $(5)$ Multiple Sampling Base Architecture.
On top of all of these architectures, we additionally evaluated each version using 1 and 2 layer encoders and decoders, as well as using KL annealing.
Figure 3: Performance metrics in 0% training dataset.
5.3 Training Hyperparameters
For training the Anomal-E dataset we used the hyperparameters present in the GitHub’s repository. These were verified to be the same hyperparameters as those mentioned in the original paper [Caville_2022].
For AutoGraphAD, we have run multiple versions of our architecture. We have run different versions with different upgrades to evaluate which yields the best results. A summary of the different settings we used can be found in Table I.
Hyperparameter Hyperparameter Values
No. Layers [1, 2]
No. Hidden 32
Learning Rate 1e−3
Activation Function ReLU
Loss Function for Structural Loss Binary Cross Entropy
Loss Function for Feature Loss [Mean Square Error, Cosine Embedding Loss]
Optimiser AdamW
Weight Decay 1e-5
KL Annealing [Yes, No]
KL Annealing Epochs 10
KL Min Weight 0.0
Regulariser [Yes, No]
Multiple Draw [Yes, No]
Maximum Epochs 100
Early Stop Patience 20
Feature Importance 1.0
Structural Importance 1.0
TABLE I: AutoGraphAD Hyperparameters
5.4 Anomaly Detection Hyperparameters
For Anomal-E to work for anomaly detection, it requires a downstream detector to characterise attacks as anomalies or not. In the case of Anomal-E we chose to use the same algorithms as they used in their original paper [Caville_2022]. These algorithms are PCA, CBLOF, and HBOS. For the Hyperparameter search, we used a more extensive search than what was used in the original paper; a summary of the Hyperparameter search can be seen in Table II.
Hyperparameter Hyperparameter Values
PCA no. components [0.96, 0.98, 0.99]
PCA Weighted True
PCA Whiten [True, False]
PCA Standardisation True
HBOS Bins [5, 6, 8, 10, 12, 14, 16, 18, 20]
HBOS Alpha [0.05, 0.1]
HBOS Tol [0.1, 0.5]
Contamination (Anomal-E) [0.02, 0.035, 0.05, 0.1, 0.2]
Alpha (AutoGraphAD) [0.1, 0.5, 1.0]
Beta (AutoGraphAD) [0.1, 0.5, 1.0]
Gamma (AutoGraphAD) [0.1, 0.5, 1.0]
MSE use (AutoGraphAD) [True, False]
Percentile (AutoGraphAD) [95, 97, 98, 99]
TABLE II: Anomal-E Estimator and AutoGraphAD Anomaly Score Hyperparameters
For AutoGraphAD, we have tried different settings to find the combination that yields the most reliable anomaly scores. As mentioned in Section 4, we have individually configurable weights that affect the anomaly score. A detailed summary of the different parameters that we have evaluated can be found in Table II.
Figure 4: Performance metrics at 3.36% contamination in the training dataset.
TABLE III: Results of Anomal-E and AutoGraphAD
0% Contamination 3.36% Contamination 5.76% Contamination
Metrics Accuracy F1 Recall Accuracy F1 Recall Accuracy F1 Recall
PCA (Anomal-E) 96.65% 82.39% 98.27% 97.3% 82,56% 95.81% 97.13% 84.24% 92.95%
CBLOF (Anomal-E) 96.21% 79.96% 94.46% 95.11% 77.51% 97.47% 97.63% 86.05% 97.76%
HBOS (Anomal-E) 96.68% 82.5% 98.28% 97.37% 84.73% 96.61% 97.1% 82.48% 91.97%
Anomaly Score (AutoGraphAD) 97.69% 84.23% 87.98% 97.67% 84.03% 87.67% 97.6% 83.36% 86.42%
TABLE IV: Time measurement of the Anomal-E’s downstream estimators and AutoGraphAD. All times are the average from all the contamination levels and are the time per graph.
Estimator Training (s) Inference (s) Anomaly Score calculation (s) Threshold calculation (s) Threshold inference (s)
PCA [pyod_pca_module] 0.0885 0.0286
CBLOF [he2003cblof] 0.0454 0.0323
HBOS [goldstein2012hbos] 0.1359 0.0869
AutoGraphAD 0.0060 0.0046 0.00529 0.000015 0.000002
5.5 Results and interpretation
For the results, we have chosen the best of our variants for each level of contamination and different weights for the anomaly score. The model hyperparameters, anomaly score weights can be found in Tables VII and VIII in the Appendix. The metrics used are Accuracy, F1 Macro and Recall Macro. These metrics were chosen as an accurate descriptor of the model’s performance and because Macro treats each class as equal [du2023skewsensitiveevaluationframeworkimbalanced], whereas other types of results do not, making them ideal for evaluating the model in a highly unbalanced dataset like UNSW-Nb15 [7348942].
Figure 5: Performance metrics of all the approaches at 5.76% contamination.
Figures 3, 4, and 5 show the results of our experimental evaluation for the different levels of contamination. AutoGraphAD yields very good results, comparable to those of Anomal-E. It should be noted that AutoGraphAD does not rely on any downstream anomaly detection algorithms such as Anomal-E, making it ideal for real-life deployment, as it can be tuned and adjusted more easily rather than requiring training on the new dataset. Furthermore, it should be noted that Anomal-E uses embeddings of size 256 [Caville_2022] whereas we rely on embeddings of size 32, 32 for the mean, and 32 for the Log Variance.
Our results at 0% contamination exceed Anomal-E’s, especially when it comes to accuracy and F1 results. Anomal-E seems to be performing better in Recall, but looking at the F1 score its Precision seems to be lower than AutoGraphAD, indicating a higher false positive count than our own approach. This continues in the case of the 3.36% contamination level, where only CBLOF seems to perform better than AutoGraphAD. In the case of 5.76% contamination levels, again we achieve higher results with the exception of Recall against most of Anomal-E’s estimators with the exception of HBOS.
It is worth noting that the size of our embeddings is eight times smaller than those used by Anomal-E. This reduction is especially important for methods that use a downstream detector, because smaller embeddings can significantly improve both performance and processing speeds [doi:10.1126/science.153.3731.34].
Figure 6: Timing for inference and training in Seconds at 0% contamination in the training dataset.
Figure 7: The timings for training and inference at 3.36% contamination.
AutoGraphAD yields substantially better inference time as seen in Table IV and Figures 6, 7, and 8. This can mainly be attributed to the fact that AutoGraphAD is GPU native, as it does not require a downstream detector as Anonmal-E, allowing for faster parallelisation and less time and resources spent on memory transfers to the CPU Anomal-E’s estimators.
Figure 8: The timings for training and inference at 5.76% contamination.
Furthermore, AutoGraphAD is capable of being more easily tuned by just adjusting the weight parameters of the anomaly score, making alterations on the anomaly score generation on the fly. Unfortunately, this cannot truly be done in the case of the estimators as their characteristics are set during initialisation. Some of the estimator’ thresholds may be manually adjusted [zhao2019pyod] [scikit-learn], but this is also the case in AutoGraphAD. Finally, in the case of model retuning or retraining, Anomal-E’s estimators need to be re-trained from the start as the PyOD’s architecture does not support retuning [zhao2019pyod]. Instead, in our case, it is very easy to retune both the model and the anomaly score, since attacks are determined from a percentile threshold that is relatively lightweight in terms of inference and training. It should be noted that the use of Robust Scaling for each sliding window’s errors allows the model to be more resilient to anomalies, since extreme values do not dominate scaling [https://doi.org/10.1002/widm.1236]. Finally, traditional anomaly detection methods are not the best when it comes to streaming environments such as NIDS [app13106353], whereas AutoGraphAD scales each graph on its own and its thresholds and anomaly score weights can be easily retuned on the fly. This makes it an ideal candidate for an NIDS system.
6 Conclusion
In conclusion, AutoGraphAD leverages recent advancements in the space of GVAEs and GAEs [GraphMAE, tian2023heterogeneousgraphmaskedautoencoders, kipf2016variationalgraphautoencoders, pan2019adversariallyregularizedgraphautoencoder] to build a more practical approach for network anomaly detection. Our design is relatively simple and yields results comparable to those of the Anomal-E. The main advantage of AutoGraphAD is faster inference and detection times, lower number of false positives, and easier and faster pipeline retuning than Anomal-E’s approach.
Furthermore, our evaluation allowed us to further explore the importance of node masking, negative edge sampling, and edge dropping. Our results highlight that lower masking, edge sampling, and edge dropping work better when we have very little or substantial contamination on our training graphs, whereas they seem more important when contamination is the same as with the rest of the dataset splits. This provides us with more information on how to tune our models for different training and inference conditions.
Our future work will focus on evaluating the generalisation capabilities of our model across unseen network scenarios, including a broader range of datasets. Assessing the capacity of a model to operate in a different environment than those seen in training is also essential to ensure practical deployment in operational networks.
Acknowledgements
This work was supported by Grant PCI2023-145974-2 funded by MICIU/AEI/10.13039/501100011033 and co-funded by the European Union (GRAPHS4SEC project).
GitHub Copilot was used to help debug and develop the code. It served mainly as a search engine for what was causing the bugs and as a recommendation engine to check which libraries and their methods are the most suitable and optimal during development.
TABLE V: NetFlow Dataset Feature Set
# Column name Feature Value NF (v9)
1 src ip Source IP address Str Yes
2 src port Source Port Int Yes
3 dst ip Destination IP address Str Yes
4 protocol Destination port Int Yes
5 ip version IP version Int Yes
6 bidirectional first seen ms Timestamp in milliseconds of the first seen packet of the flow Int Yes
7 bidirectional last seen ms Timestamp in milliseconds of the last seen packet of the flow Int Yes
8 bidirectional duration ms Duration of the entire flow in milliseconds Int Yes
9 bidirectional packets Total packets of the flow Int Yes
10 bidirectional bytes Total bytes of the flow Int Yes
11 bidirectional min packet size Minimum packet size of the flow Int Yes
12 bidirectional max packet size Maximum packet size of the flow Int Yes
13 bidirectional mean packet size Mean of the size of the packet flows Float Yes
14 bidirectional mean packet iat ms Mean of the inter arrival packet time of all the packets of the flow in milliseconds Float Yes
15 bidirectional cumulative flags Logical OR of the flags of all the packets of the flow Int Yes
16 src2dst first seen ms Timestamp in milliseconds of the first seen packet from source to destination IP address Int Yes
17 src2dst last seen ms Timestamp in milliseconds of the last seen packet from source to destination IP address Int Yes
18 src2dst duration ms Total duration of the source to destination packets in milliseconds Int Yes
19 src2dst packets Total packets sent from source to destination IP address Int Yes
20 src2dst bytes Total bytes sent from source to destination IP address Int Yes
21 src2dst min packet size Minimum packet size sent from source to destination IP address Int Yes
22 src2dst max packet size Maximum packet size sent from source to destination IP address Int Yes
23 src2dst mean packet size Mean of the size of the packets sent from source to destination IP address Float Yes
24 src2dst mean packet iat ms Mean of the inter arrival packet time of all the packets sent from source to destination IP address in milliseconds Float Yes
25 src2dst cumulative flags Logical OR of the flags of all the packets sent from source to destination IP address Int Yes
26 dst2src first seen ms Timestamp in milliseconds of the first seen packet from destination to source IP address Int Yes
27 dst2src last seen ms Timestamp in milliseconds of the last seen packet from destination to source IP address Int Yes
28 dst2src duration ms Total duration of the destination to source packets in milliseconds Int Yes
29 dst2src packets Total packets sent from destination to source IP address Int Yes
30 dst2src bytes Total bytes sent from destination to source IP address Int Yes
31 dst2src min packet size Minimum packet size sent from destination to source IP address Int Yes
32 dst2src max packet size Maximum packet size sent from destination to source IP address Int Yes
33 dst2src mean packet size Mean of the size of the packets sent from destination to source IP address Float Yes
34 dst2src mean packet iat ms Mean of the inter arrival packet time of all the packets sent from destination to source IP address in milliseconds Float Yes
35 dst2src cumulative flags Logical OR of the flags of all the packets sent from destination to source IP address Int Yes
36 classification Labeling of the dataset (0 or 1) Int No
37 category Category of the classification if present Str No
TABLE VI: Anomal-E Estimator Model Hyperparameters
Estimator Contamination Level Hyperparameters
PCA No Contamination Contamination: 0.02, Number of Components: 0.96, Whiten: False
PCA 3.5% Contamination Contamination: 0.05, Number of Components: 0.98, Whiten: False
PCA 5.7% Contamination Contamination: 0.1, Number of Components: 0.98, Whiten: False
CBLOF No Contamination Contamination: 0.02, Alpha: 0.9, Beta: 5, Number of Clusters: 36, Use Weights: True
CBLOF 3.5% Contamination Contamination: 0.05, Alpha: 0.9, Beta: 5, Number of Clusters: 40, Use Weights: True
CBLOF 5.7% Contamination Contamination: 0.1, Alpha: 0.9, Beta: 5, Number of Clusters: 50, Use Weights: False
HBOS No Contamination Contamination: 0.02, Alpha: 0.05, Beta: 5, Number of Bins: 14, Tol: 0.1
HBOS 3.5% Contamination Contamination: 0.05, Alpha: 0.1, Beta: 5, Number of Bins: 5, Tol: 0.1
HBOS 5.7% Contamination Contamination: 0.1, Alpha: 0.1, Beta: 5, Number of Bins: 12, Tol: 0.1
Figure 9: Our Anomaly Score and classification pipeline. The pipeline very closes resembles the training of our model until the first part of the Anomaly Score generation, as once we have the losses, we scale them and combine them to generate the Anomaly Scores of each connection node. Once we have each node score, we pass them through a quantile search to calculate the necessary threshold that will from then on be used to determine if a connection is anomalous or not.
TABLE VII: VGAE Model Hyperparameters
Model Variant Contamination Model Hyperparameters
VGAE Regulariser 0% 20% Negative Sampling, Edge Dropping, Node Masking with Annealing, 1 Layer
VGAE Regulariser 3.5% 40% Negative Sampling, Edge Dropping, Node Masking with Annealing, 2 Layers
VGAE Multiple Draw 5.7% 20% Negative Sampling, Edge Dropping, Node Masking with Annealing, 1 Layer, 10 Draws
TABLE VIII: VGAE Anomaly Score Hyperparameters
Contamination Anomaly Score Hyperparameters
0% Contamination Alpha: 0.1, Beta: 0.5, Gamma: 0.1, MSE: True, Percentile: 95
3.5% Contamination Alpha: 1.0, Beta: 0.1, Gamma: 1.0, MSE: False, Percentile: 95
5.7% Contamination Alpha: 0.5, Beta: 0.1, Gamma: 0.5, MSE: True, Percentile: 95

### v2

@@
-AutoGraphAD: A novel approach using Variational Graph Autoencoders for anomalous network flow detection
+AutoGraphAD: Unsupervised network anomaly detection using Variational Graph Autoencoders
 Abstract
@@
 Moreover, existing public datasets have limited and/or outdated attacks, and many of them suffer from mislabelled data.
-To reduce the reliance on labelled data, we propose AutoGraphAD, a novel unsupervised anomaly detection approach based on a Heterogeneous Variational Graph Autoencoder.
-AutoGraphAD operates on heterogeneous graphs, made from connection and IP nodes that capture network activity within a time window.
+To reduce the reliance on labelled data, we propose AutoGraphAD, a novel unsupervised anomaly detection based on a Heterogeneous Variational Graph Autoencoder.
+AutoGraphAD operates on heterogeneous graphs, made from connection and IP nodes that represent network activity.
 The model is trained using unsupervised and contrastive learning, without relying on any labelled data.
-The reconstruction, structural loss, and KL divergence are then weighted and combined in an anomaly score that is then used for anomaly detection.
-Overall, AutoGraphAD yields the same, and in some cases better, results than previous unsupervised approaches, such as Anomal-E, but without requiring costly downstream anomaly detectors.
+The model’s losses are then weighted and combined in an anomaly score used for anomaly detection.
+Overall, AutoGraphAD yields the same, and in some cases better, results than Anomal-E, but without requiring costly downstream anomaly detectors.
 As a result, AutoGraphAD achieves around 1.18 orders of magnitude faster training and 1.03 orders of magnitude faster inference, which represents a significant advantage for operational deployment.
-1 Introduction
+I Introduction
 In recent years, attacks and intrusions have been a growing problem.
@@
 A substantial body of research has explored the use of Machine Learning (ML) to detect and characterise attacks [8171733, LightweightIDS, 10127959, 10.1007/978-981-13-1921-1_35].
-However, most of these approaches are susceptible to adversarial attacks, and therefore do not suit real-life deployment [7467366].
-Graph Neural Networks are known to be more robust against adversarial attacks, since they do not rely solely on network data, but also learn the structures and relationships among network traffic flows [pujolperich2021unveilingpotentialgraphneural, 10123384, 10.1145/3664476.3664515].
+However, most of these approaches are susceptible to adversarial attacks and therefore do not suit real-life deployment [7467366].
+Graph Neural Networks (GNNs) are known to be more robust against adversarial attacks, as they use both network data and learn the structures and relationships among network traffic flows [pujolperich2021unveilingpotentialgraphneural, 10123384, 10.1145/3664476.3664515].
 In the NIDS domain, the number of datasets that can be used to create and train ML models remains limited [goldschmidt2025networkintrusiondatasetssurvey].
-Most of these datasets use outdated attacks and are highly synthetic, creating a distorted image of how attacks are structured and work [PINTO2025111177].
+Many public NIDS datasets contain synthetic elements, outdated attacks, or labelling limitations, creating a distorted image of how attacks are structured and work [PINTO2025111177].
 Labelling datasets is a costly and laborious task, which explains the scarcity of high-quality labelled data sets [Braun_2024].
@@
 Anomaly detection is a good approach to mitigate reliance on labelled datasets, as we only need normal data to establish a baseline, and anything that deviates from it is considered anomalous.
-In the domain of graph-based anomaly detection for NIDS, Anomal-E is arguably the most representative and influential approach [Caville_2022].
-Anomal-E constructs a homogeneous graph, where network flows are represented as edge features, which are then transformed to edge embeddings for anomaly detection.
-However, Anomal-E requires the generation of large embeddings, introducing significant overhead and slowing down the detection process [doi:10.1126/science.153.3731.34].
-Furthermore, in the case of concept drift [Webb2016-vo], the entire pipeline has to be re-trained, making it computationally expensive to deploy in operational environments.
-In the case of Anomal-E, and many other approaches [arganids, zoubir2024integratinggraphneuralnetworks, 10.1007/978-3-030-50423-6_12], they rely on traditional downstream anomaly detection algorithms, which are often not suitable for a streaming environment [app13106353], which makes them very difficult to deploy in real-world networks.
-In this paper, we propose AutogGraphAD, a novel approach for unsupervised anomaly detection based on Variational Graph Autoencoders (VGAEs).
+Anomal-E is arguably the most representative and influential approach in graph-based network anomaly detection [Caville_2022].
+In the case of Anomal-E and other approaches [zoubir2024integratinggraphneuralnetworks, 10.1007/978-3-030-50423-6_12], they rely on traditional downstream anomaly detection algorithms, which are often not suitable for a streaming environment [app13106353], making them very difficult to implement in real-world networks.
+This is an additional layer of computation that needs to be retuned often to account for concept drift [Webb2016-vo] and the network’s dynamic nature [app13106353].
+As both the model and the estimator would need to be taken offline to retrain and then redeploy.
+Moreover, certain one class classification algorithms tend to become prohibitively expensive when depending on the size of the embedding used.
+This can be seen in the case of OneClassSVM [NIPS1999_8725fb77] and its linear variants, where the dimensionality of the features negatively affects training and inference time, memory complexity, and overall performance [scikit-learn].
+Anomal-E [Caville_2022] uses 256 vector embeddings, which are large and can negatively affect detection algorithms.
+An interesting approach to avoid reliance on a downstream estimator is the use of dynamic Graph Autoencoders (GAEs) to detect anomalies.
+This approach has been used successfully in network monitoring as seen in GAT-AD [LATIFMARTINEZ2025110830], where the deviation of the predicted values from the actual values is used to detect anomalies.
+In this paper, we propose AutoGraphAD, a novel approach to unsupervised anomaly detection based on Variational Graph Autoencoders (VGAEs).
 In our approach, we represent network traffic as a heterogeneous graph with two distinct types of nodes, IP and Connection nodes.
 We then performed unsupervised contrastive training for the VGAE, without requiring any labelled data.
-We leverage the generated embeddings to calculate anomaly scores that are then used for anomaly detection.
+We rely solely on reconstruction errors and metrics to calculate the anomaly score, and we do not rely on any other downstream detectors.
 We have tested multiple different architectural variations to determine which implementation yields the best results at different levels of contamination (i.e., presence of anomalous flows) of the training dataset.
+The choice of testing at different contamination levels arises as we want to evaluate how robust the proposed architecture is.
 AutoGraphAD yields the same, and in some cases better, results than Anomal-E while not relying on the use of costly downstream estimators.
 Thus, faster training by 1.18 orders of magnitude and faster inference by 1.03 orders of magnitude is achieved, showcasing a significant advantage for operational deployment.
-2 Problem Description
-As mentioned in Section 1, the main problem that this paper aims to solve is the detection of network attacks without the need for expensive and limited labelled data.
-This is a very important problem that needs to be solved if we want to achieve more efficient results and a better real-life deployment of Machine Learning based NIDS.
-The main issue with current solutions such as Anomal-E [Caville_2022] and ARGANIDS [arganids] is the reliance on downstream detectors such as HBOS [goldstein2012hbos] to classify the embeddings generated for classification purposes.
-This adds an additional layer of computation that needs to be retuned often to account for concept drift [Webb2016-vo] and the network’s dynamic nature [app13106353].
-Moreover, certain one class classification algorithms tend to become prohibitively expensive when depending on the size of the embedding used.
-This can be seen in the case of OneClassSVM [NIPS1999_8725fb77] and its linear variants, where the dimensionality of the features can negatively impact training and inference time, as well as memory complexity and overall performance [scikit-learn].
-It should be noted that, in the case of Anomal-E [Caville_2022], the reported results use 256 vector embeddings.
-These embeddings can be considered large and affect anomaly detection algorithms as they suffer from the curse of dimensionality [doi:10.1126/science.153.3731.34].
-An interesting approach to avoid reliance on a downstream estimator is the use of dynamic Graph Autoencoders (VGAEs) to detect anomalies.
-This approach has been used successfully in network monitoring as seen in GAT-AD [LATIFMARTINEZ2025110830].
-In the case of GAT-AD, the predicted and actual values are used to calculate an anomaly score, and then, based on those node scores, we can detect which nodes are anomalous and which are not.
-In the case of anomaly detection in a network setting, the majority of state-of-the-art approaches rely on downstream estimators to conduct the anomaly detection.
-In this paper, our aim is to achieve network detection without relying on a downstream detector.
-We rely solely on reconstruction errors and metrics and not on any other downstream detectors.
-We are aiming to leverage VGAEs and their reconstruction errors as a way to detect anomalous nodes.
-AutoGraphAD is substantially faster computationally than Anomal-E’s and ARGANID’s approaches, while much simpler and much more easily tunable.
-3 Background
-In this section, we describe the main pre-requisites of the technologies that we will be using in this paper as part of developing our solution.
-3.1 Variational Graph Autoencoders
-The underlying technology that we will use in this paper is VGAEs. VGAEs were originally proposed by Kipf in 2016 [kipf2016variationalgraphautoencoders].
-They are a Graph Neural Network extension of traditional Variational Autoencoders [kingma2022autoencodingvariationalbayes].
-Their initial design relied on the reconstruction of the graph structure combined with the loss of Kullback-Leibler regularisation [kipf2016variationalgraphautoencoders].
-The variational autoencoder works by the encoder component creating uncertainty-aware node embeddings that are then parametrised into node embeddings and coupled with a Dot Product to recreate the Adjacency matrix of the graph [kipf2016variationalgraphautoencoders].
-Then both adjacency matrices are used to compute the reconstruction error.
-The equations used to calculate the errors can be seen in the following: Equation 1 is the Dot Product, Equation 2 is the Binary Cross Entropy Loss used to compute the structural reconstruction error [kipf2016variationalgraphautoencoders], of Equation 3 is the KL regularisation error, and Equation 4 is the total loss used.
-$A^{\prime}=Z\cdot Z^{T}$ (1)
-$L_{Recon}=BCE(A^{\prime},A)$ (2)
-$KL_{Reg}=-\frac{1}{2}*\sum(1+\log(\sigma^{2})-\mu^{2}-\sigma^{2})$ (3)
-$L_{Total}=L_{Rrecon}+KL_{Reg}$ (4)
-3.2 GraphSAGE
-GraphSAGE is a Graph Neural Network architecture that can learn to generate embeddings from sampling and aggregating the node’s neighbours [NIPS2017_5dd9db5e].
-After the aggregation, the generated embedding goes through a non-linear update.
-The main advantage of GraphSAGE compared to other algorithms is that it can easily work with unseen nodes [NIPS2017_5dd9db5e], making it ideal for real-time deployment where the nodes and graph structure differ from training data.
-In addition, GraphSAGE has been proven to scale very well in larger graphs [NIPS2017_5dd9db5e], making it an ideal choice for a large network deployment.
-3.3 Adversarially Regularised VGAEs
-Adversarially Regularised VGAEs [pan2019adversariallyregularizedgraphautoencoder] is another approach to train the autoencoders to generate higher quality embeddings.
-This approach uses the exact same approach as the original VGAE paper [kipf2016variationalgraphautoencoders] but adds the use of a regulariser model to improve the performance.
-The regulariser works by taking samples from the generated embeddings and from the a Gaussian distribution and learns to separate which one is which.
-This is then fed into the total loss that is used for back-propagation.
-3.4 Multiple Sampling Variational Autoencoders
-Another upgrade that can be done to enhance the VGAEs is through the use of multiple latent embeddings per node.
-The idea behind using multiple embeddings that are then used to reconstruct multiple reconstruction losses.
-This allows for the generation of an unbiased estimate of the same objective with a much smaller variance [pmlr-v48-mnihb16].
-This has been proven to improve the performance of the model and yield better results [pmlr-v80-buchholz18a].
-3.5 KL Annealing
-KL Annealing can be used to improve the performance of VAEs [bowman2016generating].
-KL Annealing works by slowly increasing the KL regularisation loss through a set number of epochs.
-By slowly increasing the effect of the KL loss, the autoencoder is initially forced to learn meaningful embeddings, and then transitions on enforcing prior matching.
-The goal of this approach is to mitigate the posterior collapse, where KL transitions to 0, which can negatively affect the performance of the VAE.
-3.6 Graph Masked Autoencoders
-Masking was proposed by Hou et al.
-[GraphMAE] in the Graph Masked Autoencoder (GraphMAE) and used in the Heterogenous Graph Masked Autoencoder (HGMAE) [tian2023heterogeneousgraphmaskedautoencoders] to enhance the performance of their model.
-Their GAE architecture is enhanced by masking the input node features as well as the latent embeddings.
-Additionally, as part of model training, the graph edges are dropped.
-For masking, they use a learnable masking token similar to the one used in BERT [devlin2019bertpretrainingdeepbidirectional].
-Masking and edge dropping make it harder for the autoencoder to reconstruct the information, and thus the autoencoder does not learn trivial solutions during training but rather focusses on learning the node relations.
-4 Methodology
-In this section, we focus on describing our methodology and how we have created, trained, and evaluated AutoGraphAD.
-Additionally, we describe our pre-processing, graph-building, and dataset creation pipelines.
-4.1 Dataset Used
-For this work, we have chosen to use the UNSW-NB15 dataset [7348942].
+We evaluate the method under multiple contamination levels and show comparable detection quality to Anomal-E while substantially reducing training and inference time.
+II Related Work
+NIDS are a broad area with thousands of proposals; a review of all proposals is outside the scope of this paper.
+Thus, we will focus on related work in terms of GNN-based NIDS.
+Surveys on NIDS and GNN-based NIDS can be found here [goldschmidt2025networkintrusiondatasetssurvey, ZHONG2024103821].
+GNN-based NIDS are mostly supervised.
+E-GraphSAGE [Lo_2022] is a widely used representative baseline as it is a widely evaluated flow-graph GNN [Venturi2023].
+The model works by implementing E-GraphSAGE, a variation of GraphSAGE that works by processing only the Edge features rather than the node features that traditional GraphSAGE captures [NIPS2017_5dd9db5e].
+It is trained to perform edge (flow) classification.
+GraphIDS is a recently published model based on E-GraphSAGE [guerra2025selfsupervisedlearninggraphrepresentations] by augmenting flow representation learning with a masked encoder-decoder Transformer.
+ARGANIDS [arganids] uses Adversarially Regularised Graph Autoencoder (ARGA) [pan2019adversariallyregularizedgraphautoencoder] GAE to learn node embeddings and then uses a downstream supervised Machine Learning classifier.
+As mentioned in the Introduction, the main limitation of using supervised learning is the reliance on limited and possibly outdated datasets [PINTO2025111177] due to the increased cost of labelling network data [Braun_2024].
+This cannot easily scale for a real-life implementation, leading to the need for unsupervised approaches.
+When it comes to unsupervised approaches, there is much less related work.
+Anomal-E is a state of the art network anomaly detection algorithm that implements a GAE to learn meaningful embeddings of a graph [Caville_2022].
+Anomal-E implements E-GraphSAGE.
+A modified Graph Infomax version is then used to train the model encoder.
+The resulting embeddings are then fed to a downstream estimator used for anomaly detection.
+Although this approach is effective, the pipeline adds more complexity through additional components and tuning options.
+As such, we compare primarily unsupervised approaches, as we assume that there are no labels available during training.
+III Methodology
+In this section, we first describe the dataset construction and graph representation.
+We then focus on describing our methodology and how we have created, trained, and evaluated AutoGraphAD.
+Finally, we explain the anomaly scoring and evaluation.
+III-A Dataset
+For this work, we used the UNSW-NB15 dataset [7348942].
 We chose UNSW-NB15 for the AutoGraphAD evaluation, as it is a widely used dataset with clearly labelled modern attacks.
 Thus, our evaluation is better suited for the current environment.
-The data set has been generated from the raw PCAP files using NFStream [AOUINI2022108719].
-For the generation of the flows, we used the same settings as were used in the original dataset generation.
-NFStream can output flows with multiple features.
-We chose the features that overlap with the features present in NetFlow V9 [claise2004cisco].
-For UNSW-NB15, we used the generated labels from the original paper [7348942] and matched them with the flows generated by NFStream. The features used can be found in Table V.
-4.2 Dataset Creation
-Having described the data pre-processing, it is important to describe how the datasets are created.
+The data set has been generated from the raw PCAP files using NFStream [AOUINI2022108719], which outputs flows with NetFlow V9 compatible features [claise2004cisco].
+To create the labels, we used the UNSW-NB15 generated labels [7348942] and matched them to the generated flows.
 The goal of this paper is to explore and showcase our model’s ability to detect anomalies even when the training data are contaminated.
 To achieve this, we have created different training parts with different levels of contamination similar to Anomal-E [Caville_2022].
-The main difference between Anomal-E’s approach and ours is that Anomal-E uses 10% of the data set to create a single large graph for the training, and the same is repeated for testing.
-Our goal is to create 3 levels of contamination: No contamination, 3.5% Contamination, and 5.7% Contamination.
-In our case, we use time windows to detect attacks and simulate a real-life deployment.
-For the training split of the dataset, we follow the following procedure; once we create all the sliding windows for the entire dataset, we calculate if the windows are anomalous or not, meaning if they contain attacks or not, and we do a stratified splitting to all datasets.
-For a training dataset that does not have any attacks, we removed all the anomalous graphs.
-For 3.5% contamination, we keep the stratified split, and for 5.7% we downsample the benign graphs for UNSW-NB15 as the normal contamination of the level is around 3.5%.
-4.3 Feature Pre-processing and Graph Creation
-For feature pre-processing, we use a similar pre-processing approach to the one proposed in Anomal-E [Caville_2022].
-The dataset is initially checked for flows with missing or infinite values, which are then removed.
-Following this step, we begin by building the time windows that will be used for training, testing, and evaluation.
-We then split the randomly generated graphs into parts for training, evaluation, and testing.
+The main difference between Anomal-E’s approach and ours is that Anomal-E uses 10% of the data set to create a single large graph for training, and the same is repeated for testing.
+We use time windows to detect attacks and simulate real-life deployments.
+To achieve a fair distribution of attacks, we use stratified data split to see whether or not contaminated attacks are equally distributed.
+For 0% contamination, we remove all the graphs with attacks, for contamination, we use a stratified split of the dataset, and for 5.7% we downsample the benign graphs to of their original number.
 We allocate 70% of the data for training, 20% for evaluation, and 10% for testing.
 For the time windows, we chose to create non-overlapping graphs that capture 180 seconds of traffic.
-The second step of the pre-processing is feature normalisation.
-For categorical features, which in our case is the IP version and protocols, we employ One-Hot encoding [Hancock2020].
-In the case of the protocols, we use Rare Labelling from the library Feature Engine [Galli2021] due to the incredibly large number of protocols and the fact that the majority of these are very sparse in the data.
-Including these features in One-Hot encoding would increase the dimensionality of the data without yielding any performance gain; as such, the rarer protocols are bundled in a third category labelled Rare, resulting in three One-Hot categories, two for the two most common features, and one for all the rare features.
-For the continuous features, we are using L2 row wise normalisation.
-This choice was made because it does not require any significant pre-processing and can be done on the fly in a real-life environment, allowing for an easier real-life deployment.
-The entire process from dataset creation to graph creation can be seen in Figure 1.
-Figure 1: Dataset Pre-processing pipeline for graph generation
-4.4 Network Architecture and Variants
+For categorical features, which in our case is the IP version and protocols, we employ One-Hot encoding.
+In the case of the protocols, we use Rare Labelling from the library Feature Engine [Galli2021] to reduce the dimensionality of the data.
+For continuous features, we use L2 row-wise normalisation.
+The graph is made up of IP nodes that contain placeholder values used for message passing and connection nodes that contain flow features. The graph is undirected.
 Figure 1: The proposed Architecture of AutoGraphAD in a training setting.
@@
 The pipeline starts with the encoder that generates the latent space embeddings that are then reparametised to create the embeddings that are then used for structure and feature reconstruction.
-Using the reconstructed values, we calculate the reconstruction losses used for back-propagation.
-The GNN encoder and decoder can be switched out with every other GNN algorithm.
-The same can be done regarding the losses that are used in Back-propagation.
-Our proposed architecture builds upon the original VGAE [kipf2016variationalgraphautoencoders] and extends it by incorporating elements from GraphMAE [GraphMAE] and HGMAE [tian2023heterogeneousgraphmaskedautoencoders].
-Our architecture is designed to reconstruct both node features and graph structure during training.
+The GNN encoder and decoder as well as the losses can be switched out with every other GNN algorithm.
+In this paper, we are employing GraphSAGE as the encoder and decoder algorithm.
+III-B Model Architecture
+Our proposed architecture builds on the original VGAE [kipf2016variationalgraphautoencoders] and extends it by incorporating elements from GraphMAE [GraphMAE] and HGMAE [tian2023heterogeneousgraphmaskedautoencoders].
+Our architecture is designed to reconstruct both the node features and the graph structure during training.
 The main upgrades introduced in our design are as follows.
-We are using the same masking component as used in GraphMAE [GraphMAE] and HGMAE [tian2023heterogeneousgraphmaskedautoencoders], but instead of masking both latent embeddings and original graph embeddings, we only mask the original graph nodes.
+We are using the same masking component as used in GraphMAE [GraphMAE] and HGMAE [tian2023heterogeneousgraphmaskedautoencoders], but we only mask the original graph nodes.
 We made this decision because during the evaluation of AutoGraphAD there was a negligible improvement when using double masking.
@@
 We have also incorporated negative edge sampling as a contrastive learning objective to help the autoencoder to better distinguish between real and fake edges.
-This helps our autoencoder to learn better the structure of the graphs.
 For structural reconstruction, we use an enhanced Dot Product.
@@
 $A^{\prime}=Z\cdot(W*Z^{T})$ (1)
-Regarding feature reconstruction, a GNN has been employed to decode the node features.
+Regarding feature reconstruction, a GNN model has been employed to decode only the connection node features.
 A GNN was chosen as it is the most capable of decoding the embeddings correctly.
 GNNs as a decoder are used in both HGMAE and GraphMAE [tian2023heterogeneousgraphmaskedautoencoders, GraphMAE], but in these approaches they have focused only on decoding masked embeddings.
-In this architecture, all nodes are decoded.
-In our approach, we only decode the connection nodes as the IP nodes are placeholder values that do not contain any meaningful information.
+In this architecture, all of the connection nodes are decoded.
 Regarding the feature reconstruction errors, we have created two main variants, one using the Cosine Embedding Error and one using the Mean Squared Error (MSE).
-The Cosine Embedding Error approach has been inspired by HGMAE [tian2023heterogeneousgraphmaskedautoencoders] and GraphMAE [GraphMAE], while the MSE approach has been chosen as the goal here of the autoencoder is to be able to generate meaningful embeddings in which the decoded values must be closest to the actual values.
+The Cosine Embedding Error approach has been inspired by HGMAE [tian2023heterogeneousgraphmaskedautoencoders] and GraphMAE [GraphMAE], while the MSE approach was chosen to help generate meaningful embeddings that are closest to the actual values.
 The KL divergence loss used in our model is the same as in the original VGAE paper [kipf2016variationalgraphautoencoders].
 For the total reconstruction loss, we combine the feature reconstruction loss with the structural reconstruction loss and KL divergence.
-The total loss is then back-propagated through the model to update the model.
-Each component of the total reconstruction loss, with the exception of the KL divergence, is weighted by a user defined value.
+Each component of the total reconstruction loss, with the exception of the KL divergence, is weighted by a user-defined value.
 The model architecture can be found in Figure 1 and the total loss is depicted in Equation 2.
-$L_{Total}=\alpha*L_{Structure}+\beta*L_{Features}+KL$ (6)
-4.5 Model Training
+$L_{Total}=\alpha*L_{Struct}+\beta*L_{Feat}+KL$ (2)
 The model was trained with 100 maximum allowable epochs.
@@
 Furthermore, we saved the model’s weights at the point that the model achieved its best performance.
-We ensured that we maintain the optimal version of the model.
-4.6 Anomaly Detection
-For anomaly detection, we are calculating the anomaly scores of the connection nodes.
+III-C Anomaly Detection
+For anomaly detection, we calculate the anomaly scores of the connection nodes.
 The anomaly scores are based on the calculation of the reconstruction errors with the KL term.
 The idea is that the anomaly score will be lower for normal nodes and higher for anomalous nodes.
-To determine the threshold of what an anomaly is and what it is not, we use percentiles to determine a user defined percentile, and then everything above it is determined as anomalous scores.
+We use a percentile-based threshold, where a node is classified as anomalous if its score exceeds the p-th percentile; $p$ is selected offline as a threshold hyperparameter.
 For our score, we use the following equation.
 $Score=\alpha*L_{Feat}+\beta*L_{Struct}+\gamma*KL$ (3)
-The equation works by using a user-assigned score to weight the importance of each loss in the anomaly score.
-This allows us to use different weights for different settings.
-Additionally, we employ Robust Scaling [https://doi.org/10.1002/widm.1236] of all reconstruction errors and KL Divergence that are used for the anomaly scores.
-This guarantees that all scores are on the same scale and that no value dominates the range and skews the anomaly scores [10.1145/2594473.2594476].
+The equation works by using user-assigned weights to weigh the importance of each loss in the anomaly score.
+This allows us to use different weights for different settings and adapt to the changing nature of the network.
+We employ Robust Scaling [https://doi.org/10.1002/widm.1236] of all reconstruction errors and KL Divergence that are used for anomaly scores to ensure that no value dominates the scale.
 Our entire approach runs in PyTorch and on the GPU.
-Providing a faster processing time that using Sci-kit learn [scikit-learn].
-The pipeline can be found in Figure 9 in the Appendix.
-4.7 Evaluating Inference Performance
-To evaluate the performance of each method, we have chosen to measure the training and inference times.
-For timing, we used the Python built-in library, time.
-More specifically, we used the function perf_counter_ns() which gives us the most accurate measurement for our experiment, since we are measuring short durations.
-For timing, we do not count the parts where data transfer takes place to make the comparisons more equal.
-Moreover, in the case of AutoGraphAD, we can time each individual pass when generating the Anomaly Score, but this is not possible during threshold setting and threshold inference timing and when timing Anomal-E’s downstream estimator training and inference.
-For these cases, we get the total time and divide it based on the number of graphs that were processed.
-For the recorded timings, in the case of AutoGraphAD we add the average time per graph during the anomaly score with its respective time in the threshold component.
-For example, in the case of the training time, we add the average anomaly score generation time per graph with the estimated threshold calculation time per graph.
-4.8 Libraries Used
-To create our VGAE, we used PyTorch [NEURIPS2019_bdbca288] and an extension of PyTorch called PyTorch Geometric [fey2019fastgraphrepresentationlearning] which is best suited for the creation and training of GNNs.
-To ensure easier reproducibility and faster training, we employed PyTorch Lightning [Falcon_PyTorch_Lightning_2019].
-Regarding anomaly detection, we employed NumPy [harris2020array] and PyOD [zhao2019pyod].
-Finally, for data pre-processing, window creation, and running the downstream tasks for Anomal-E, we used Scikit-learn [scikit-learn] and Pandas [mckinney-proc-scipy-2010].
-To run the Anomal-E code we used DGL [wang2020deepgraphlibrarygraphcentric].
-5 Experimental Results
-5.1 Experimental setup
-To evaluate AutoGraphAD, we used UNSW-NB15 in both AutoGraphAD and Anomal-E. For Anomal-E, we used an implementation found in a GitHub repository [^1] https://github.com/waimorris/Anomal-E.
-Overall, we trained the models in the different contamination levels [^2] The contamination level indicates the percentage of anomalous flows present in the training set.
-Ideally, anomaly detection methods should be trained only on normal flows.
-However, this is difficult to guarantee in real-world environments.
-Hence, robustness to a small degree of contamination is desirable for practical deployment in operational networks.
-followed by the anomaly detection estimators’ grid search for Anomal-E and weight search for the Anomaly Score in the VGAE implementation.
+III-D Evaluating Inference Performance
+To evaluate the performance of each method, we have chosen to measure the training and inference times. For timing, we used the Python built-in library, time.
+We do not count the parts where data transfer takes place to make the comparisons more equal.
+Moreover, in the case of AutoGraphAD, we can time each individual pass by generating the Anomaly Score.
+However, we cannot time each pass in the threshold setting, threshold inference timing, and Anomal-E’s downstream estimator training and inference timing.
+In that form, we get the average pass time by dividing the total by all of the passes.
+III-E Libraries Used
+For AutoGraphAD, we used PyTorch, PyTorch Geometric [fey2019fastgraphrepresentationlearning] and DGL [wang2020deepgraphlibrarygraphcentric] for the GNNs.
+For faster training, we employed PyTorch Lightning [Falcon_PyTorch_Lightning_2019].
+NumPy and PyOD [zhao2019pyod] are used for anomaly detection.
+For data pre-processing, window creation, and running downstream tasks for Anomal-E, we used Scikit-learn [scikit-learn] and Pandas.
+IV Experimental Results
+IV-A Experimental setup
+To evaluate AutoGraphAD, we used UNSW-NB15 in both AutoGraphAD and Anomal-E.
+For Anomal-E, we used an implementation found in a GitHub repository [^1] https://github.com/waimorris/Anomal-E.
+We chose to only evaluate Anomal-E as it is a pure Anomaly Detection paper.
+Overall, we trained the models in the different contamination levels followed by the anomaly detection estimators’ grid search for Anomal-E and weight search for the Anomaly Score in the VGAE implementation.
 During the hyperparameter search, the model weights were frozen.
+Labels are used only for offline evaluation and score-weight selection for anomaly detection.
 To run our experiments, we used a server with an Nvidia RTX 3090 with 24 GB of VRAM.
@@
 The server came with 64 GB of RAM and its operating system was Ubuntu 22.04.4 LTS.
-5.2 VGAE Variants
+IV-B VGAE Variants
 We exhaustively tested multiple different VGAE variants.
 Our base architecture is node masking and edge dropping to force the variational autoencoders to learn non-trivial node relationships.
-The variants that we evaluated are the following: $(1)$ Base architecture using MSE for node feature reconstruction, $(2)$ Base architecture using Cosine Embedding Error for node feature reconstruction, $(3)$ Regularised Base Architecture, $(4)$ Split Base Architecture, and $(5)$ Multiple Sampling Base Architecture.
-On top of all of these architectures, we additionally evaluated each version using 1 and 2 layer encoders and decoders, as well as using KL annealing.
+The variants that we evaluated are the following: $(1)$ Base architecture using MSE for node feature reconstruction, $(2)$ Base architecture using Cosine Embedding Error for node feature reconstruction, $(3)$ Regularised Base Architecture, $(4)$ Split Base Architecture and $(5)$ Multiple Sampling Base Architecture.
+We evaluated each version using 1 and 2 layer encoders and decoders, as well as using KL annealing. The implementation of all our variants are available in the accompanying code repository.
 Figure 2: Performance metrics in 0% training dataset.
-5.3 Training Hyperparameters
+IV-C Training Hyperparameters
 For training the Anomal-E dataset we used the hyperparameters present in the GitHub’s repository.
@@
 We have run different versions with different upgrades to evaluate which yields the best results.
-A summary of the different settings we used can be found in Table I.
-Hyperparameter Hyperparameter Values
-No. Layers [1, 2]
-No. Hidden 32
-Learning Rate 1e−3
-Activation Function ReLU
-Loss Function for Structural Loss Binary Cross Entropy
-Loss Function for Feature Loss [Mean Square Error, Cosine Embedding Loss]
-Optimiser AdamW
-Weight Decay 1e-5
-KL Annealing [Yes, No]
-KL Annealing Epochs 10
-KL Min Weight 0.0
-Regulariser [Yes, No]
-Multiple Draw [Yes, No]
-Maximum Epochs 100
-Early Stop Patience 20
-Feature Importance 1.0
-Structural Importance 1.0
-TABLE I: AutoGraphAD Hyperparameters
-5.4 Anomaly Detection Hyperparameters
-For Anomal-E to work for anomaly detection, it requires a downstream detector to characterise attacks as anomalies or not.
+We relied on Edge Index accuracy to select the best version of our model.
+IV-D Anomaly Detection Hyperparameters
+Anomal-E requires a downstream detector to characterise attacks as anomalies or not.
 In the case of Anomal-E we chose to use the same algorithms as they used in their original paper [Caville_2022].
-These algorithms are PCA, CBLOF, and HBOS.
-For the Hyperparameter search, we used a more extensive search than what was used in the original paper; a summary of the Hyperparameter search can be seen in Table II.
-Hyperparameter Hyperparameter Values
-PCA no. components [0.96, 0.98, 0.99]
-PCA Weighted True
-PCA Whiten [True, False]
-PCA Standardisation True
-HBOS Bins [5, 6, 8, 10, 12, 14, 16, 18, 20]
-HBOS Alpha [0.05, 0.1]
-HBOS Tol [0.1, 0.5]
-Contamination (Anomal-E) [0.02, 0.035, 0.05, 0.1, 0.2]
-Alpha (AutoGraphAD) [0.1, 0.5, 1.0]
-Beta (AutoGraphAD) [0.1, 0.5, 1.0]
-Gamma (AutoGraphAD) [0.1, 0.5, 1.0]
-MSE use (AutoGraphAD) [True, False]
-Percentile (AutoGraphAD) [95, 97, 98, 99]
-TABLE II: Anomal-E Estimator and AutoGraphAD Anomaly Score Hyperparameters
+For the Hyperparameter search, we used a more extensive search space, the grid can be found in our repository [^2] https://github.com/georgeani/AutoGraphAD/.
+Our Anomal-E implementation, including all variants evaluated, is available in our code repository.
+We used a Brute Force Hyperparameter search where we used the dataset labels to tune the parameters.
+We chose the results that yielded the best balance between F1-Score, Precision, and Recall.
+Figure 3: Performance metrics at 3.36% contamination in the training dataset.
+TABLE I: Results of Anomal-E and AutoGraphAD
 For AutoGraphAD, we have tried different settings to find the combination that yields the most reliable anomaly scores.
-As mentioned in Section 4, we have individually configurable weights that affect the anomaly score.
-A detailed summary of the different parameters that we have evaluated can be found in Table II.
-Figure 4: Performance metrics at 3.36% contamination in the training dataset.
-TABLE III: Results of Anomal-E and AutoGraphAD
+A detailed summary of the different parameters that we have evaluated can be found in our repository.
 0% Contamination 3.36% Contamination 5.76% Contamination
@@
 Anomaly Score (AutoGraphAD) 97.69% 84.23% 87.98% 97.67% 84.03% 87.67% 97.6% 83.36% 86.42%
-TABLE IV: Time measurement of the Anomal-E’s downstream estimators and AutoGraphAD. All times are the average from all the contamination levels and are the time per graph.
-Estimator Training (s) Inference (s) Anomaly Score calculation (s) Threshold calculation (s) Threshold inference (s)
-PCA [pyod_pca_module] 0.0885 0.0286
-CBLOF [he2003cblof] 0.0454 0.0323
-HBOS [goldstein2012hbos] 0.1359 0.0869
-AutoGraphAD 0.0060 0.0046 0.00529 0.000015 0.000002
-5.5 Results and interpretation
+IV-E Results and interpretation
 For the results, we have chosen the best of our variants for each level of contamination and different weights for the anomaly score.
-The model hyperparameters, anomaly score weights can be found in Tables VII and VIII in the Appendix.
 The metrics used are Accuracy, F1 Macro and Recall Macro.
-These metrics were chosen as an accurate descriptor of the model’s performance and because Macro treats each class as equal [du2023skewsensitiveevaluationframeworkimbalanced], whereas other types of results do not, making them ideal for evaluating the model in a highly unbalanced dataset like UNSW-Nb15 [7348942].
+These metrics were chosen as an accurate descriptor of the model’s performance and because Macro treats each class as equal [du2023skewsensitiveevaluationframeworkimbalanced], making them ideal for evaluating the model in a highly unbalanced dataset like UNSW-NB15 [7348942].
+At 0% and 3.36% contamination, we used our regularised variant, and at 5.76% we used the multiple sampling variant.
 Figure 4: Performance metrics of all the approaches at 5.76% contamination.
-Figures 3, 4, and 5 show the results of our experimental evaluation for the different levels of contamination.
+Figures 2, 3, and 4 show the results of our experimental evaluation for the different levels of contamination.
 AutoGraphAD yields very good results, comparable to those of Anomal-E.
 It should be noted that AutoGraphAD does not rely on any downstream anomaly detection algorithms such as Anomal-E, making it ideal for real-life deployment, as it can be tuned and adjusted more easily rather than requiring training on the new dataset.
-Furthermore, it should be noted that Anomal-E uses embeddings of size 256 [Caville_2022] whereas we rely on embeddings of size 32, 32 for the mean, and 32 for the Log Variance.
 Our results at 0% contamination exceed Anomal-E’s, especially when it comes to accuracy and F1 results.
 Anomal-E seems to be performing better in Recall, but looking at the F1 score its Precision seems to be lower than AutoGraphAD, indicating a higher false positive count than our own approach.
-This continues in the case of the 3.36% contamination level, where only CBLOF seems to perform better than AutoGraphAD.
-In the case of 5.76% contamination levels, again we achieve higher results with the exception of Recall against most of Anomal-E’s estimators with the exception of HBOS.
-It is worth noting that the size of our embeddings is eight times smaller than those used by Anomal-E.
-This reduction is especially important for methods that use a downstream detector, because smaller embeddings can significantly improve both performance and processing speeds [doi:10.1126/science.153.3731.34].
-Figure 6: Timing for inference and training in Seconds at 0% contamination in the training dataset.
-Figure 7: The timings for training and inference at 3.36% contamination.
-AutoGraphAD yields substantially better inference time as seen in Table IV and Figures 6, 7, and 8.
-This can mainly be attributed to the fact that AutoGraphAD is GPU native, as it does not require a downstream detector as Anonmal-E, allowing for faster parallelisation and less time and resources spent on memory transfers to the CPU Anomal-E’s estimators.
-Figure 8: The timings for training and inference at 5.76% contamination.
-Furthermore, AutoGraphAD is capable of being more easily tuned by just adjusting the weight parameters of the anomaly score, making alterations on the anomaly score generation on the fly.
+This continues at the 3.36% contamination level, where only CBLOF seems to perform better than AutoGraphAD.
+In the case of 5.76% contamination levels, again we achieve higher results with the exception of Recall against most of Anomal-E’s estimators.
+AutoGraphAD yields substantially better inference time (0.009892 s) compared to PCA (0.0286 s), CBLOF (0.0323 s) and HBOS(0.0869 s).
+This can also be observed in training: AutoGraphAD (0.011305 s), PCA (0.0885 s), CBLOF(0.0454 s), HBOS (0.1359 s).
+This can mainly be attributed to the fact that AutoGraphAD does not require a downstream detector as Anomal-E.
+AutoGraphAD is capable of being more easily tuned by simply adjusting the weight parameters of the anomaly score, making alterations to the anomaly score generation on the fly.
 Unfortunately, this cannot truly be done in the case of the estimators as their characteristics are set during initialisation.
-Some of the estimator’ thresholds may be manually adjusted [zhao2019pyod] [scikit-learn], but this is also the case in AutoGraphAD.
+Some of the estimator’s thresholds may be manually adjusted [zhao2019pyod] [scikit-learn], but this is also the case in AutoGraphAD.
 Finally, in the case of model retuning or retraining, Anomal-E’s estimators need to be re-trained from the start as the PyOD’s architecture does not support retuning [zhao2019pyod].
-Instead, in our case, it is very easy to retune both the model and the anomaly score, since attacks are determined from a percentile threshold that is relatively lightweight in terms of inference and training.
-It should be noted that the use of Robust Scaling for each sliding window’s errors allows the model to be more resilient to anomalies, since extreme values do not dominate scaling [https://doi.org/10.1002/widm.1236].
-Finally, traditional anomaly detection methods are not the best when it comes to streaming environments such as NIDS [app13106353], whereas AutoGraphAD scales each graph on its own and its thresholds and anomaly score weights can be easily retuned on the fly.
-This makes it an ideal candidate for an NIDS system.
-6 Conclusion
+Instead, AutoGraphAD is very easy to retune both the model and the anomaly score due to its design.
+V Conclusion
 In conclusion, AutoGraphAD leverages recent advancements in the space of GVAEs and GAEs [GraphMAE, tian2023heterogeneousgraphmaskedautoencoders, kipf2016variationalgraphautoencoders, pan2019adversariallyregularizedgraphautoencoder] to build a more practical approach for network anomaly detection.
 Our design is relatively simple and yields results comparable to those of the Anomal-E.
-The main advantage of AutoGraphAD is faster inference and detection times, lower number of false positives, and easier and faster pipeline retuning than Anomal-E’s approach.
-Furthermore, our evaluation allowed us to further explore the importance of node masking, negative edge sampling, and edge dropping.
-Our results highlight that lower masking, edge sampling, and edge dropping work better when we have very little or substantial contamination on our training graphs, whereas they seem more important when contamination is the same as with the rest of the dataset splits.
-This provides us with more information on how to tune our models for different training and inference conditions.
+The main advantage of AutoGraphAD is faster inference and detection times, a lower number of false positives, and easier and faster pipeline retuning than Anomal-E’s approach.
 Our future work will focus on evaluating the generalisation capabilities of our model across unseen network scenarios, including a broader range of datasets.
@@
 Acknowledgements
-This work was supported by Grant PCI2023-145974-2 funded by MICIU/AEI/10.13039/501100011033 and co-funded by the European Union (GRAPHS4SEC project).
+This work was supported by Grant PCI2023-145974-2 funded by MICIU/AEI/10.13039/501100011033 and cofunded by the European Union (GRAPHS4SEC project).
+This work is also supported by the Catalan Institution for Research and Advanced Studies (ICREA Academia).
 GitHub Copilot was used to help debug and develop the code.
 It served mainly as a search engine for what was causing the bugs and as a recommendation engine to check which libraries and their methods are the most suitable and optimal during development.
-TABLE V: NetFlow Dataset Feature Set
-# Column name Feature Value NF (v9)
-1 src ip Source IP address Str Yes
-2 src port Source Port Int Yes
-3 dst ip Destination IP address Str Yes
-4 protocol Destination port Int Yes
-5 ip version IP version Int Yes
-6 bidirectional first seen ms Timestamp in milliseconds of the first seen packet of the flow Int Yes
-7 bidirectional last seen ms Timestamp in milliseconds of the last seen packet of the flow Int Yes
-8 bidirectional duration ms Duration of the entire flow in milliseconds Int Yes
-9 bidirectional packets Total packets of the flow Int Yes
-10 bidirectional bytes Total bytes of the flow Int Yes
-11 bidirectional min packet size Minimum packet size of the flow Int Yes
-12 bidirectional max packet size Maximum packet size of the flow Int Yes
-13 bidirectional mean packet size Mean of the size of the packet flows Float Yes
-14 bidirectional mean packet iat ms Mean of the inter arrival packet time of all the packets of the flow in milliseconds Float Yes
-15 bidirectional cumulative flags Logical OR of the flags of all the packets of the flow Int Yes
-16 src2dst first seen ms Timestamp in milliseconds of the first seen packet from source to destination IP address Int Yes
-17 src2dst last seen ms Timestamp in milliseconds of the last seen packet from source to destination IP address Int Yes
-18 src2dst duration ms Total duration of the source to destination packets in milliseconds Int Yes
-19 src2dst packets Total packets sent from source to destination IP address Int Yes
-20 src2dst bytes Total bytes sent from source to destination IP address Int Yes
-21 src2dst min packet size Minimum packet size sent from source to destination IP address Int Yes
-22 src2dst max packet size Maximum packet size sent from source to destination IP address Int Yes
-23 src2dst mean packet size Mean of the size of the packets sent from source to destination IP address Float Yes
-24 src2dst mean packet iat ms Mean of the inter arrival packet time of all the packets sent from source to destination IP address in milliseconds Float Yes
-25 src2dst cumulative flags Logical OR of the flags of all the packets sent from source to destination IP address Int Yes
-26 dst2src first seen ms Timestamp in milliseconds of the first seen packet from destination to source IP address Int Yes
-27 dst2src last seen ms Timestamp in milliseconds of the last seen packet from destination to source IP address Int Yes
-28 dst2src duration ms Total duration of the destination to source packets in milliseconds Int Yes
-29 dst2src packets Total packets sent from destination to source IP address Int Yes
-30 dst2src bytes Total bytes sent from destination to source IP address Int Yes
-31 dst2src min packet size Minimum packet size sent from destination to source IP address Int Yes
-32 dst2src max packet size Maximum packet size sent from destination to source IP address Int Yes
-33 dst2src mean packet size Mean of the size of the packets sent from destination to source IP address Float Yes
-34 dst2src mean packet iat ms Mean of the inter arrival packet time of all the packets sent from destination to source IP address in milliseconds Float Yes
-35 dst2src cumulative flags Logical OR of the flags of all the packets sent from destination to source IP address Int Yes
-36 classification Labeling of the dataset (0 or 1) Int No
-37 category Category of the classification if present Str No
-TABLE VI: Anomal-E Estimator Model Hyperparameters
-Estimator Contamination Level Hyperparameters
-PCA No Contamination Contamination: 0.02, Number of Components: 0.96, Whiten: False
-PCA 3.5% Contamination Contamination: 0.05, Number of Components: 0.98, Whiten: False
-PCA 5.7% Contamination Contamination: 0.1, Number of Components: 0.98, Whiten: False
-CBLOF No Contamination Contamination: 0.02, Alpha: 0.9, Beta: 5, Number of Clusters: 36, Use Weights: True
-CBLOF 3.5% Contamination Contamination: 0.05, Alpha: 0.9, Beta: 5, Number of Clusters: 40, Use Weights: True
-CBLOF 5.7% Contamination Contamination: 0.1, Alpha: 0.9, Beta: 5, Number of Clusters: 50, Use Weights: False
-HBOS No Contamination Contamination: 0.02, Alpha: 0.05, Beta: 5, Number of Bins: 14, Tol: 0.1
-HBOS 3.5% Contamination Contamination: 0.05, Alpha: 0.1, Beta: 5, Number of Bins: 5, Tol: 0.1
-HBOS 5.7% Contamination Contamination: 0.1, Alpha: 0.1, Beta: 5, Number of Bins: 12, Tol: 0.1
-Figure 9: Our Anomaly Score and classification pipeline.
-The pipeline very closes resembles the training of our model until the first part of the Anomaly Score generation, as once we have the losses, we scale them and combine them to generate the Anomaly Scores of each connection node.
-Once we have each node score, we pass them through a quantile search to calculate the necessary threshold that will from then on be used to determine if a connection is anomalous or not.
-TABLE VII: VGAE Model Hyperparameters
-Model Variant Contamination Model Hyperparameters
-VGAE Regulariser 0% 20% Negative Sampling, Edge Dropping, Node Masking with Annealing, 1 Layer
-VGAE Regulariser 3.5% 40% Negative Sampling, Edge Dropping, Node Masking with Annealing, 2 Layers
-VGAE Multiple Draw 5.7% 20% Negative Sampling, Edge Dropping, Node Masking with Annealing, 1 Layer, 10 Draws
-TABLE VIII: VGAE Anomaly Score Hyperparameters
-Contamination Anomaly Score Hyperparameters
-0% Contamination Alpha: 0.1, Beta: 0.5, Gamma: 0.1, MSE: True, Percentile: 95
-3.5% Contamination Alpha: 1.0, Beta: 0.1, Gamma: 1.0, MSE: False, Percentile: 95
-5.7% Contamination Alpha: 0.5, Beta: 0.1, Gamma: 0.5, MSE: True, Percentile: 95